OpenStudy (anonymous):

1/sqrt x-6 what is the domain and range

OpenStudy (anonymous):

I know the domain is (6,infinity) i can't figure out if the range is (0,infinity) or (-infinity,infinity)

OpenStudy (anonymous):

\[\frac{ 1 }{ \sqrt{x-6} }\]

OpenStudy (anonymous):

@shamil98

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@abb0t

OpenStudy (anonymous):

@marsss

OpenStudy (anonymous):

@marsss ? where'd you go

OpenStudy (anonymous):

for the domain, there must be a positive expression inside square so x-6>0 x>6 for the range, y=1/sqrt(x-6) 0<y<1 f: (6,∞) -> (0,1]